OpenStudy (anonymous):

A boy is 6 ft. tall. The distance from the boy to a mirror is 8 ft. From the mirror to the house is 16 ft. How high is the top of the house?

OpenStudy (anonymous):

I think this one needs a picture....

OpenStudy (anonymous):

this one didn't come with any picture. :( @AnimalAin

OpenStudy (anonymous):

Unless there is some more information, I can't help you. Sorry.

OpenStudy (anonymous):

Hold on I think I can solve this.

OpenStudy (anonymous):

6/8 = h/16 Solve that and tell me what you got.

OpenStudy (anonymous):

6/8 = h/16 8h = 96 h=12

OpenStudy (anonymous):

Is this the correct answer? @vzfreakz

OpenStudy (anonymous):

Yes! Good job :)